# Non-affine Families of 8 x 8 Complex Hadamard Matrices

### v1

@@ 2.1 Families of complex Hadamard matrices
The general theory and applications of Hadamard matrices are covered in [1] and [12]. The complex case has been studied more comprehensively in [2, 19, 22].
@@ 2.2 Palindromic polynomials
Notice that the discriminant of (2.1) is the discriminant of (2.2) scaled with $x_{2}^{\alpha_{2}}\cdots x^{\alpha_{n}}_{n}|p_{2}|$. In practice it can be more convenient to consider the inequality $C/(x_{2}^{\alpha_{2}}\cdots x^{\alpha_{n}}_{n})\leq 0$ instead of $|p_{1}|/|p_{2}|\leq 2$, where $C$ is the discriminant of (2.1). Moreover, Lemma 2.2 and Lemma 2.3 can be shown to work for anti-palindromic polynomials as well with almost identical proofs.
@@ 3.1 Affine families
$\begin{bmatrix}1&1&1&1&1&1&1&1\\ 1&\omega&\omega^{2}&\omega^{3}&\omega^{4}&\omega^{5}&\omega^{6}&\omega^{7}\\ 1&\omega^{2}&\omega^{6}&\omega^{4}&1&\omega^{2}&\omega^{6}&\omega^{4}\\ 1&\omega^{3}&\omega^{6}&\omega&\omega^{4}&\omega^{7}&\omega^{2}&\omega^{5}\\ 1&\omega^{4}&\omega^{4}&1&1&\omega^{4}&\omega^{4}&1\\ 1&\omega^{5}&\omega^{2}&\omega^{7}&\omega^{4}&\omega&\omega^{6}&\omega^{3}\\ 1&\omega^{6}&\omega^{2}&\omega^{4}&1&\omega^{6}&\omega^{2}&\omega^{4}\\ 1&\omega^{7}&\omega^{6}&\omega^{5}&\omega^{4}&\omega^{3}&\omega^{2}&\omega,\end{bmatrix}\circ\begin{bmatrix}1&1&1&1&1&1&1&1\\ 1&a&b&c&1&a&b&c\\ 1&d&1&d&1&d&1&d\\ 1&e&b&\overline{a}ce&1&e&b&\overline{a}ce\\ 1&1&1&1&1&1&1&1\\ 1&a&b&c&1&a&b&c\\ 1&d&1&d&1&d&1&d\\ 1&e&b&\overline{a}ce&1&e&b&\overline{a}ce\end{bmatrix}$
@@ 4.1 Family $T_{8C}^{(3)}$
It can be noticed that if $(a,b,c,d,e,f)$ is a solution to the system in Lemma 4.1, then so is $(a,b,d^{2}/c,d,e,f)$. We also have that $\mathscr{D}(T_{8C})^{T}(a,b,c,d,e,f)=\mathscr{D}(T_{8C})(a,b,d^{2}/c,d,e)$. However, it does not always hold that $\mathscr{D}(\mathcal{S}_{C})^{T}(a,b,c)=\mathscr{D}(\mathcal{S}_{C})(a,b,(d^{\prime})^{2}/c)$, which would imply that $T_{8C}^{(3)}$ is symmetric, as the solution $(a,b,d^{2}/c,d,e,f)$ can also be in one of the other solution branches. On the other hand, as the different branches seem to give equivalent families, we conjecture that $T^{(3)}_{8C}$ is symmetric.
@@ 4.4 Family $T^{(3)}_{8F}$
We can solve the system (4.1) by introducing different variables for real and imaginary parts and adding unimodularity constraints between them.
@@
$T_{8C}^{(3)}$ is a 3-paremeter family of complex Hadamard matrices.
@@ 5 Inequivalence
Suppose that $f(x_{1},\ldots,x_{k})$ is an iteratively invertible function with iterable entries $\phi_{1},\ldots\phi_{k}$ at indices $(i_{1},j_{1}),\ldots,(i_{k},j_{k})$. We can now compute $(x_{1}^{\prime},\ldots,x^{\prime}_{k})$ from $H=f(x_{1}^{\prime},\ldots,x^{\prime}_{k})$ with
$(\phi^{-1}_{1}(h_{i_{1},j_{1}}),\;\phi^{-1}_{2}(\phi^{-1}_{1}(h_{i_{1},j_{1}}),h_{i_{2},j_{2}}),\ldots,\phi^{-1}_{k}(\phi_{1}^{-1}(h_{i_{1},j_{1}}),\ldots,h_{i_{k},j_{k}})),$ (5.1)
@@
    For each $H^{\prime}=(h^{\prime}_{i,j})$, perform a backtracking search to find parameters $x^{\prime}=(x^{\prime}_{1},\dots,x^{\prime}_{k})$, such that $\langle f_{i,j}(x^{\prime})\rangle=\langle h^{\prime}_{i,j}\rangle$, by evaluating the formula (5.1) with elements from $\{h^{\prime}_{i,j}\}$.
@@ 5.2 Inequivalence of the families
$\begin{split}F^{(5)}_{8},\;D^{(5)}_{8A},\;D^{(5)}_{8B},\;&S_{8A}^{(4)},\;S^{(4)}_{8B},\\ T_{8B}^{(3)},\;T_{8C}^{(3)},\;T^{(3)}_{8D},\;(T^{(3)}_{8D})^{T},&\;T^{(3)}_{8E},\;(T^{(3)}_{8E})^{T},\;T^{(3)}_{8F}.\end{split}$ (5.2)
Let $f:D\to\mathbb{H}(8)$ be the function defining $H^{(k)}_{8}$ as given in Sections 3 and 4. Consider $H^{\prime}=f(e^{2i},e^{3i},\dots,e^{(k+1)i})\in\mathbb{H}(8)$. The only family from the list (5.2) to which $H^{\prime}$ belongs is $H_{8}^{(k)}$ itself.
@@
One can check that the point $(e^{2i},e^{3i},\dots,e^{(k+1)i})$ belongs to the domains of all the families in the list (5.2). An Exhaustive computer search with the algorithm described in Section 5.1 yields the result. ∎
An immediate corollary of Theorem 5.1 is that all the families in the list (5.2) are pairwise inequivalent, and no family is contained within another. One can also confirm that the point $(e^{2i},e^{3i},\ldots,e^{ki})$ is not a boundary point of any of the domains, meaning there exists a neighbourhood around that point belonging uniquely to the corresponding family.
@@ 5.3 Butson-type matrices
In [21], all BH$(8,4)$ Butson-type matrices are placed in some affine family from the list (3.1). For BH$(8,6)$ matrices, such a categorisation is not possible. Consider a BH$(8,6)$ matrix
@@
and its inequivalent transpose $B_{2}=B_{1}^{T}$, where $\omega=\exp(2\pi i/8).$
@@
The matrices $B_{1}$ and $B_{2}$ do not belong in any of the families in the list (5.2).

### v2

@@ 2.1 Families of complex Hadamard matrices
The general theory and applications of Hadamard matrices are covered in [1] and [11]. The complex case has been studied more comprehensively in [18, 21].
@@ 2.2 Palindromic polynomials
Notice that the discriminant of (1) is the discriminant of (2) scaled with $x_{2}^{\alpha_{2}}\cdots x^{\alpha_{n}}_{n}|p_{2}|$. In practice it can be more convenient to consider the inequality $C/(x_{2}^{\alpha_{2}}\cdots x^{\alpha_{n}}_{n})\leq 0$ instead of $|p_{1}|/|p_{2}|\leq 2$, where $C$ is the discriminant of (1). Moreover, Lemma 2.5 and Lemma 2.6 can be shown to work for anti-palindromic polynomials as well with almost identical proofs.
@@ 3.1 Affine families
$\begin{bmatrix}1&1&1&1&1&1&1&1\\ 1&\omega&\omega^{2}&\omega^{3}&\omega^{4}&\omega^{5}&\omega^{6}&\omega^{7}\\ 1&\omega^{2}&\omega^{6}&\omega^{4}&1&\omega^{2}&\omega^{6}&\omega^{4}\\ 1&\omega^{3}&\omega^{6}&\omega&\omega^{4}&\omega^{7}&\omega^{2}&\omega^{5}\\ 1&\omega^{4}&\omega^{4}&1&1&\omega^{4}&\omega^{4}&1\\ 1&\omega^{5}&\omega^{2}&\omega^{7}&\omega^{4}&\omega&\omega^{6}&\omega^{3}\\ 1&\omega^{6}&\omega^{2}&\omega^{4}&1&\omega^{6}&\omega^{2}&\omega^{4}\\ 1&\omega^{7}&\omega^{6}&\omega^{5}&\omega^{4}&\omega^{3}&\omega^{2}&\omega,\end{bmatrix}\circ\begin{bmatrix}1&1&1&1&1&1&1&1\\ 1&a&b&c&1&a&b&c\\ 1&d&d&1&1&d&d&1\\ 1&e&b&\overline{a}ce&1&e&b&\overline{a}ce\\ 1&1&1&1&1&1&1&1\\ 1&a&b&c&1&a&b&c\\ 1&d&d&1&1&d&d&1\\ 1&e&b&\overline{a}ce&1&e&b&\overline{a}ce\end{bmatrix}$
@@ 4.1 Family $T_{8C}^{(3)}$
It can be noticed that if $(a,b,c,d,e,f)$ is a solution to the system in Lemma 4.1, then so is $(a,b,d^{2}/c,d,e,f)$. We also have that
$\mathscr{D}(T_{8C})^{T}(a,b,c,d,e,f)=\mathscr{D}(T_{8C})(a,b,d^{2}/c,d,e).$
However, it does not always hold that $\mathscr{D}(\mathcal{S}_{C})^{T}(a,b,c)=\mathscr{D}(\mathcal{S}_{C})(a,b,(d^{\prime})^{2}/c)$, which would imply that $T_{8C}^{(3)}$ is symmetric, as the solution $(a,b,d^{2}/c,d,e,f)$ can also be in one of the other solution branches. On the other hand, as the different branches seem to give equivalent families, we conjecture that $T^{(3)}_{8C}$ is symmetric.
@@ 4.4 Family $T^{(3)}_{8F}$
We can solve the system (7) by introducing different variables for real and imaginary parts and adding unimodularity constraints between them.
@@
$T_{8C}^{(3)}$ is a 3-parameter family of complex Hadamard matrices.
@@ 5 Inequivalence
Suppose that $f(x_{1},\ldots,x_{k})$ is an iteratively invertible function with iterable entries $\phi_{1},\ldots,\phi_{k}$ at indices $(i_{1},j_{1}),\ldots,(i_{k},j_{k})$. We can now compute $(x_{1}^{\prime},\ldots,x^{\prime}_{k})$ from $H=f(x_{1}^{\prime},\ldots,x^{\prime}_{k})$ with
$(\phi^{-1}_{1}(h_{i_{1},j_{1}}),\;\phi^{-1}_{2}(\phi^{-1}_{1}(h_{i_{1},j_{1}}),h_{i_{2},j_{2}}),\ldots,\phi^{-1}_{k}(\phi_{1}^{-1}(h_{i_{1},j_{1}}),\ldots,h_{i_{k},j_{k}})),$ (8)
@@
    For each $H^{\prime}=(h^{\prime}_{i,j})$, perform a backtracking search to find parameters $x^{\prime}=(x^{\prime}_{1},\dots,x^{\prime}_{k})$, such that $\langle f_{i,j}(x^{\prime})\rangle=\langle h^{\prime}_{i,j}\rangle$, by evaluating the formula (8) with elements from $\{h^{\prime}_{i,j}\}$.
@@ 5.2 Inequivalence of the families
$\begin{split}F^{(5)}_{8},\;D^{(5)}_{8A},\;D^{(5)}_{8B},\;&S_{8A}^{(4)},\;S^{(4)}_{8B},\\ T_{8B}^{(3)},\;T_{8C}^{(3)},\;T^{(3)}_{8D},\;(T^{(3)}_{8D})^{T},&\;T^{(3)}_{8E},\;(T^{(3)}_{8E})^{T},\;T^{(3)}_{8F}.\end{split}$ (9)
Let $f:D\to\mathbb{H}(8)$ be the function defining $H^{(k)}_{8}$ as given in Sections 3 and 4. Consider $H^{\prime}=f(e^{2i},e^{3i},\dots,e^{(k+1)i})\in\mathbb{H}(8)$. The only family from the list (9) to which $H^{\prime}$ belongs is $H_{8}^{(k)}$ itself.
@@
One can check that the point $(e^{2i},e^{3i},\dots,e^{(k+1)i})$ belongs to the domains of all the families in the list (9). An Exhaustive computer search with the algorithm described in Section 5.1 yields the result. ∎
An immediate corollary of Theorem 5.2 is that all the families in the list (9) are pairwise inequivalent, and no family is contained within another. One can also confirm that the point $(e^{2i},e^{3i},\ldots,e^{ki})$ is not a boundary point of any of the domains, meaning there exists a neighbourhood around that point belonging uniquely to the corresponding family.
@@ 5.3 Butson-type matrices
In [20], all BH$(8,4)$ Butson-type matrices are placed in some affine family from the list (3). For BH$(8,6)$ matrices, such a categorisation to known families is not possible as there are matrices that do not belong any of the families from the list (9). Consider a BH$(8,6)$ matrix
@@
and its inequivalent transpose $B_{2}=B_{1}^{T}$, where $\omega=\exp(2\pi i/6).$
@@
The matrices $B_{1}$ and $B_{2}$ do not belong in any of the families in the list (9).